OpenStudy (anonymous):

Find the area under the graph of the function over the given interval??? y=e^x;[-5,9]

OpenStudy (fibonaccichick666):

so, have you talked about riemann sums?

OpenStudy (anonymous):

Riemann sums? Slow down a bit, this is just basic integration. \[\int\limits_{-5}^{9}e^xdx\]

OpenStudy (anonymous):

can you write it down step by step please?

OpenStudy (fibonaccichick666):

you have to cover those before you are allowed to integrate

OpenStudy (fibonaccichick666):

no, you will have to work on this with us, we need to know where you are in class first

OpenStudy (anonymous):

\[ \int\limits_{-5}^{9}e^xdx = e^9 - e^5 \]

OpenStudy (anonymous):

so basically b-a

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Why on Earth would you use a Riemann sum for this?

OpenStudy (fibonaccichick666):

that is the fundamental theorem of calculus in action

OpenStudy (anonymous):

isn't it -1/e^9

OpenStudy (anonymous):

sorry -1/e^5****

OpenStudy (anonymous):

I made a mistake in my last one, it should be e^-5

OpenStudy (fibonaccichick666):

because prior to integration you have to learn riemann sums

OpenStudy (anonymous):

\[ \int\limits_{-5}^{9}e^xdx = e^9 - e^{-5} \]

OpenStudy (freckles):

I don't think I would want to do riemann sums with the exponential function involved :p

OpenStudy (fibonaccichick666):

and @kevst3r, a negative exponent doesn't yield a negative answer, but due to FTC you have a minus un front

OpenStudy (anonymous):

in other words, \[e^9-1/e^5\]

OpenStudy (fibonaccichick666):

I was asking to see if we were allowed to use plain old integration(to us)

OpenStudy (freckles):

unless we could use finite amount of rectangles

OpenStudy (fibonaccichick666):

-5 is probably better notation, but yes

OpenStudy (freckles):

i might be okay with that if the finite number is low

OpenStudy (anonymous):

Yes @kevst3r, both your way and my way of expressing it are correct but I prefer keeping it on one line.

OpenStudy (fibonaccichick666):

You have to learn sums prior to learning integration and memorizing the integrals so I just needed a reference point

OpenStudy (anonymous):

i have another question can i just ask it here?

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

or do i have to create another question

OpenStudy (anonymous):

ok nice

OpenStudy (anonymous):

i have two more questions for you guys then i'm good

OpenStudy (anonymous):

okay, \[y=(1/2)x^2, y=-x^2+6\]

OpenStudy (anonymous):

the answer is 16, but i'm not sure how you do this problem

OpenStudy (anonymous):

Well how would you integrate 0.5x^2?

OpenStudy (anonymous):

Sorry, I know this isn't relevant but can someone take a look at my question ? I really need help

OpenStudy (anonymous):

^dude get off my question go make ur own

OpenStudy (anonymous):

I did but no one ever replys this is the only way i can actually get some help

OpenStudy (anonymous):

jesus christ

OpenStudy (anonymous):

tom any advice for this problem?

OpenStudy (fibonaccichick666):

how do you integrate something of the form \[x^n\]?

OpenStudy (freckles):

just to make sure or to make a guess you are finding the area between the curves you have presented just now?

OpenStudy (anonymous):

find the area bounded by the given curves \[y=\frac{ 1 }{ 2 }(x^2),y=-x^2+6\]

OpenStudy (anonymous):

yes freckles

OpenStudy (fibonaccichick666):

ah, then sketch first

OpenStudy (anonymous):

the answer is 16 but how

OpenStudy (anonymous):

i dont remember learning this one

OpenStudy (freckles):

do you know how to find the intersections of the equations?

OpenStudy (freckles):

that would be my first step

OpenStudy (anonymous):

isn't that where you make them equal each other?

OpenStudy (anonymous):

so for example

OpenStudy (freckles):

and yea a graph would help like fibo suggested

OpenStudy (anonymous):

1/2x^2=-x^2+6

OpenStudy (anonymous):

multiply both sides

OpenStudy (freckles):

yes

OpenStudy (fibonaccichick666):

a sketch will help a lot

OpenStudy (anonymous):

i did that part, i got 2 points, -2 and 2

OpenStudy (anonymous):

but idk what to do next

OpenStudy (anonymous):

x= +2 and -2

OpenStudy (fibonaccichick666):

draw a picture.... I swear you won't regret it

OpenStudy (anonymous):

sweetie, during a final exam, i don't have time to draw a graph

OpenStudy (anonymous):

why not a lesbesque integral i'd love to see that here

OpenStudy (fibonaccichick666):

ya do, I swear

OpenStudy (freckles):

this is a final exam?

OpenStudy (anonymous):

no but i have one coming up in 2 days

OpenStudy (anonymous):

and i have this problem and another one that i need help on if i want to ace it

OpenStudy (anonymous):

so i get two points, -2 and +2, what do i do afterwards?

OpenStudy (fibonaccichick666):

draw a picture I say again, and it shall make it clear

OpenStudy (freckles):

well i like pictures but you can plug in a number between the intersections to see which function is greater

OpenStudy (anonymous):

freckles thats another thing i dont understand, how do you find the greater function

OpenStudy (anonymous):

i think it's -x^2+6?

OpenStudy (freckles):

well first off both of these equations are parabolas not that hard to draw but if you absolutely refused to draw a graph plug in a number between the intersections -2 and 2. like 0 plug in 0 into 1/2*x^2 and 0 into -x^2+6

OpenStudy (fibonaccichick666):

A picture is a fantastic way

OpenStudy (anonymous):

-x^2+6 cause you're left with 6 lol

OpenStudy (anonymous):

whats next?

OpenStudy (freckles):

integrate the difference of the larger-smaller on the interval you found

OpenStudy (anonymous):

i don't know why i didn't find out about this website earlier, i love this

OpenStudy (anonymous):

so (1/2x^2 )-(-x^2+6)?

OpenStudy (freckles):

well you said the larger was -x^2+6

OpenStudy (freckles):

on that interval -2 to 2

OpenStudy (anonymous):

sorry i ment the other way around

OpenStudy (freckles):

(-x^2+6)-(1/2x^2) integrate that on the interval -2 to 2

OpenStudy (anonymous):

(-x^2+6)-(1/2x^2)

OpenStudy (anonymous):

what do you mean integrate t hat on the intervale -2 to 2

OpenStudy (fibonaccichick666):

find the integral of (-x^2+6)-(1/2x^2) on [-2,2]

OpenStudy (anonymous):

last question guys

OpenStudy (anonymous):

this one is stupid easy but i just need confirmation

OpenStudy (anonymous):

y=23/x;y=16 when x=1

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

hellooo?

OpenStudy (anonymous):

whered my geniuses go

ganeshie8 (ganeshie8):

hey you seem new around here

ganeshie8 (ganeshie8):

let me take a moment to welcome you :) \[\color{green}{\text{Welcome to Openstudy!}}\] Please limit to one question per thread and you need to close the question to open a new question

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

do i have to open a new thread for each question

ganeshie8 (ganeshie8):

yes because the new questions show up in top of the list and your chances for getting help quick increase scroll up to the start of the question to close it